OpenStudy (anonymous):

the perimeter of a rectangle is 210inches. the length exceeds the width by 61 inches. find the length and width

OpenStudy (anonymous):

2l+2b=210 l-b=61 l=83 b=22

OpenStudy (gw2011):

I agree with the above answer. Here is how I solved the problem: Let w=width Let l=length=w+61 P=2w+2l P=2w+2(w+61) 210=2w+2w+122 210-122=2w+2w 88=4w 22=w l=w+61=22+61=83 Therefore, the width=22 and the length=83